OpenStudy (anonymous):

Consider the equation 4x + 6y = 48. Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points) Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points)

OpenStudy (anonymous):

HELP PLease ill rank you

OpenStudy (aripotta):

one sec, i'm brushing up on the different methods

OpenStudy (anonymous):

ok:) @AriPotta

OpenStudy (aripotta):

ok, well first we have to change the equation into y = mx + b

OpenStudy (aripotta):

4x + 6y = 48 6y = -4x + 48 y = -4/6 + 8

OpenStudy (anonymous):

thanks @AriPotta was that part one

OpenStudy (aripotta):

so we know that the y-intercept is 8

OpenStudy (aripotta):

no, not yet lol.

OpenStudy (aripotta):

i'm working this out, sorry

OpenStudy (aripotta):

so we put a point at (0,8)

OpenStudy (anonymous):

its fine :) youre helping its all good

OpenStudy (aripotta):

then use the slope, -4/6 to plot another point. we go down 4 and go to the right 6 to get a point at (6,5)

OpenStudy (anonymous):

tell me when we being part 2

OpenStudy (aripotta):

so with two points, we can make the line (except it's going to be straight, and not caddywonkus)

OpenStudy (aripotta):

so that was part 1

OpenStudy (aripotta):

use what i was saying for your written part

OpenStudy (aripotta):

and now part 2

OpenStudy (anonymous):

:)

OpenStudy (aripotta):

for the intercepts method, we can use our y = mx + b equation y = -4/6x + 8

OpenStudy (aripotta):

wait, i think i'm doing something wrong..

OpenStudy (aripotta):

sigh.

OpenStudy (anonymous):

hmm whats that

OpenStudy (aripotta):

the intercepts method isn't really working out for me..

OpenStudy (anonymous):

mmm iseee

OpenStudy (aripotta):

give me a few minutes to check...

OpenStudy (aripotta):

ahh i see

OpenStudy (aripotta):

ok, we have to find the x-intercept. so we substitute 0 for y in our equation. 0 = -4/6x + 8 -8 = -4/6x 12 = x so our x-intercept is 12. and our first point will be at (12,0) then we can plot the y-intercept, which is at (0,8). then we can draw the line in between them

OpenStudy (aripotta):

there you go :)

OpenStudy (aripotta):

sorry i had to extend the graph for the x-intercept lol

OpenStudy (anonymous):

thanks :)!

OpenStudy (aripotta):

does that make sense though?

OpenStudy (anonymous):

yes i actaully undeerstood it

OpenStudy (aripotta):

great :)

OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (-5, 9). y = -5 y = 9 x = -5 x = 9

OpenStudy (aripotta):

y = 9